the fence around yard is 66 meters long. the length of the yard is 12 meters, what is the width of the yard

Answer by Math_Boss(45)

p=perimeter
p=66
l=12
w=?
p=2(l+w)
66=2(12+w)
66=24+2w
42=2w
w=21
Ans:21
Check:12+12+21+21=66